### v1

@@ III RESULTS AND DISCUSSIONS
Ideally, the reference direction for $\theta$ should align with the applied electric field. However, in our system, the electrodes lie within the same plane (the xy-plane), and the effective field distribution is spatially non-uniform. Therefore, using a fixed axis direction provides a more consistent basis for orientational analysis. Although the applied electric field lies primarily along the x-direction, we analyzed the orientational order parameter with respect to the z-axis ($C_{z}$) because the water molecules are confined near the xy-plane. The z-direction captures out-of-plane dipole tilting and fluctuations more sensitively, which are critical to revealing the reorientation mechanisms under field perturbations. This approach has been commonly adopted in interfacial water studies to characterize structural disruption and ordering normal to the surface [44]. For this definition, $C=0$ corresponds to random dipole orientations, $C=1$ indicates perfect alignment along the z-axis, and $C=-\frac{1}{2}$ represents dipole alignment perpendicular to the z-axis.
@@
To understand this more deeply, the order parameter profile for fluid near the bottom plate for 4 V AC and DC voltages has been analyzed and shown in Fig. 9. In this case slabs contain water within 1 Å of the bottom plate, and in different regions along the plate. For the DC voltage, the left (blue) and right (red) electrodes correspond to negative and positive charges, respectively. The order parameter profile shows a peak on top of both electrodes, indicating that the maximum deviation from the initial ordering occurs near the electrode positions. Interestingly, the peaks are symmetrical for AC potential and asymmetrical with a large peak on the negative electrode for the DC potential. With this, we can obtain a complete picture of the water molecule ordering near the electrode surface. In the absence of an applied electric field, water molecules at the gold interface exhibit preferential orientation driven by subtle van der Waals and image charge interactions. Studies have shown that the molecular dipole of water tends to lie parallel to the gold surface, resulting in an in-plane orientation and a negative order parameter ($C_{z}<0$), as also observed in the present study [45, 46]. In this configuration, the oxygen atom (O) typically lies closer to the gold surface, while the hydrogen atoms (H) point slightly away from the surface. This geometry arises because oxygen interacts more favorably with the surface, potentially through weak hydrogen bonding or van der Waals attraction, while the hydrogen atoms avoid direct interaction with gold due to their weaker or slightly repulsive interactions. As a result, the dipole vector (from O to the midpoint of the H–H bond) lies parallel to the surface or is slightly tilted away, leading to a well-defined in-plane ordering in the interfacial water layer.
@@
When an electric potential is applied, the orientation of these dipoles changes in response to the polarity of the electrodes. At the negative electrode, the positively charged hydrogen atoms are attracted toward the surface. This causes the dipoles to tilt in the opposite direction compared to their initial configuration, disrupting the original in-plane alignment. As a result, the order parameter $C_{z}$ increases significantly toward zero, indicating a strong reorientation or near-random dipole alignment. This explains why, under DC potential, the $C_{z}$ peak on top of the negative electrode rises to nearly zero.
@@
In contrast, at the positive electrode, the oxygen atoms, which were already slightly tilted toward the surface, experience further attraction due to their negative partial charge. This reinforces the initial orientation rather than disrupting it, resulting in only a modest change in the order parameter. Consequently, the $C_{z}$ peak on the positive side remains closer to 0.18-0.18, not reaching zero. This asymmetry in reorientation between the two electrodes under DC voltage accounts for the asymmetric peak heights in the $C_{z}$ profile observed in Fig. 9. For the AC case, the field reverses direction periodically, which averages out these polarity-specific effects, leading to a more symmetrical response in the order parameter profile.
@@
Until now, we have been focussing on symmetrical electrodes with the same dimensions. It has been already reported that there is a net flow of electrolyte solutions when AC potential is applied to asymmetric pairs of the electrode structure [24, 26]. Here we focus on three different models to study the influence of asymmetry of electrodes. The first electrode structure is a symmetric pair of electrodes with both electrodes of the same length D and the distance between the electrodes is also fixed at D. The second structure is an asymmetric electrode in which the left electrode length is D and the right electrode length is 2D (twice the length of left electrode) and the distance between the electrodes is D. Similarly, the third electrode structure is an asymmetric electrode in which the left electrode length is D and the right electrode is 2D, however the distance between electrodes is 2D. These models are named as model 1, model 2, and model 3 respectively. The structure and dimensions of the three models are shown in Fig. 10. MD simulations were carried out on these three models using a constant potential method on the electrodes with a 4 V AC potential having a frequency of 100 GHz. The $x$ component of the fluid velocity in the $x$-$y$ plane for the three models is plotted in Fig. 10. The velocity profile is averaged over five different independent simulations. Due to the periodicity of the system, even in the cases where the electrodes have different lengths there is symmetry (through reflection + translation). This means that if flow is produce by the AC voltage, it is equally likely to be in either direction. Therefore, in averaging the five independent simulations, the average velocities reported are relative to the sign of the overall velocity of the first simulation. The average value of the $x$ velocity component in the entire channel is given in the legend of the plots of Fig. 10. From the velocity profile, it is evident that a net flow in the $x$ direction is obtained for asymmetric electrode configuration. Whereas the net velocity of the symmetric electrode structure is zero, within the statistical uncertainty. It is also noted that model 3 gives maximum velocity. The average value of the $x$ component of model 3 is 0.78 $\pm$ 0.34 m s-1. A parabolic flow with maximum velocity at the center of the channel is obtained. Similarly, model 2 also shows a net $x$ component velocity of 0.23 $\pm$ m 0.09s-1.
@@
Figure 10 also shows the streamline plot of three models. The streamlines are more random and chaotic in model 1. The absence of a significant net flow in any direction is also confirmed by the average velocity profile which indicates no net flow. This implies that the forces acting on the system are balanced. It results in a homogeneous distribution of the streamlines with no dominance of the directional movement.
@@
The third model displays very directional streamlines, thereby indicating a strong and coherent net flow. This suggests that the forces driving the flow are strong and dominant over turbulent influences. The mean velocity profile shows a net flow in the $x$-direction having a velocity component of 0.78 $\pm$ 0.34 m s-1. This net flow in the $x$ direction, together with the streamlined directionality pattern, suggests there is a preferential flow direction. The asymmetries in forces acting on the system drive the fluid preferentially along the $x$ axis over the period considered.
@@
The mean square displacement (MSD) were calculated to understand the fluid properties, and the results are shown in Fig. 11. These systems have been analyzed under different electric field conditions: 0.0 V (no field), 4.0 V DC , and 4.0 V at 100 GHz AC. The MSD plots help to depict how the position of the particles changes with time. At 0.0 V, a linear increase in MSD indicates normal diffusive behavior. At 4.0 V DC, the MSD is close to that of the no-field condition, indicating minimal impact on the mobility. However, the 100 GHz AC electric field of 4.0 V results in a significantly increased MSD, indicating that the mobility of the particles is being enhanced. This implies that the high-frequency field disrupts particle interactions, thus leading to more displacement over time.
@@
Our study aimed to uncover the molecular-level mechanisms responsible for the AC-EHD flow. Several explanations regarding the underlying mechanism of the AC-EHD flow have been proposed in the literature [7, 1]. Here, by combining the results of our MD simulation and previous understandings, we provide further insight into the mechanism. Various important factors are responsible for this phenomenon. First, localized heat generation occurs near the electrodes due to the effect of the rapid reorientation of water dipoles under the high-frequency AC potential [47]. This localized heating gives rise to a steep temperature gradient from the hot region near the electrodes to cooler regions far from them. This temperature disparity causes a change in water density, leading to a density gradient. This density gradient drives buoyancy-driven convection, where less dense heated water rises, and denser cooler water descends to replace it, creating convective currents that enhance the movement of fluid [48].
Secondly, electrothermal effects also play a significant role. The electric field changes the temperature-dependent properties of water, such as its conductivity and permittivity, to give rise to these effects [7]. This gradient in conductivity and permittivity, in turn, interacts with the electric field to produce electrothermal forces that drive the fluid movement. These electrothermal forces are further increased for higher frequencies, thereby enhancing the fluid movement [5].
@@
Nevertheless, a net flow is only visible when the electrodes are asymmetric. In symmetric electrode configurations, the aforementioned forces are balanced, thereby eliminating the chances of a net directional flow. The difference in the electrode dimensions and the distance between electrodes will enhance the imbalance in forces acting on the fluid, leading to a pronounced AC-EHD effect. Thus the high-frequency AC potential enhances mobility while inducing localized heating near the electrodes. The combined effect of thermal expansion, density gradients, buoyancy-driven convection, electrothermal forces, and Maxwell stress drives the flow under AC-EHD conditions. Notably, the asymmetry in electrode configuration is vital for creating an imbalance in these forces and that would result in a net flow. These insights provide a comprehensive understanding of the molecular-level mechanism of AC-EHD flow in nanoscale systems under very high frequency ranges, which are crucial for optimizing and controlling fluid flows in various applications.
@@ IV CONCLUSIONS
Several physical mechanisms were identified as contributors to this flow: (1) buoyancy-driven convection resulting from temperature-induced density gradients, (2) electrothermal forces from the interaction between the electric field and temperature-dependent properties of water, and (3) Maxwell stress and electrostriction caused by spatially varying polarization. The observed flow is enabled only when these forces are imbalanced, which occurs under geometrically asymmetric electrode configurations.
@@ Acknowledgements.
This research was supported by the Australian Research Council through the ARC Laureate Fellowship program (FL190100080). It was also supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia and resources and services from the National Computational Infrastructure (NCI) under the Adapter scheme, which is supported by the Australian Government. We also wish to acknowledge The University of Queensland’s Research Computing Centre (RCC) and Queensland Cyber Infrastructure Foundation (QCIF) for their support.

### v2

@@ III RESULTS AND DISCUSSIONS
Ideally, the reference direction for $\theta$ should align with the applied electric field. However, in our system, the electrodes lie within the same plane (the xy-plane), and the effective field distribution is spatially non-uniform. Therefore, using a fixed axis direction provides a more consistent basis for orientational analysis. Although the applied electric field lies primarily along the x-direction, we analyzed the orientational order parameter with respect to the z-axis ($C_{z}$) because the water molecules are confined near the xy-plane. The z-direction captures out-of-plane dipole tilting and fluctuations more sensitively, which are critical to revealing the reorientation mechanisms under field perturbations. This approach has been commonly adopted in interfacial water studies to characterize structural disruption and ordering normal to the surface [willard2010instantaneous]. For this definition, $C=0$ corresponds to random dipole orientations, $C=1$ indicates perfect alignment along the z-axis, and $C=-\frac{1}{2}$ represents dipole alignment perpendicular to the z-axis.
@@
To understand this more deeply, the order parameter profile for fluid near the bottom plate for 4 V AC and DC voltages has been analyzed and shown in Fig. 9. In this case slabs contain water within 1 Å of the bottom plate, and in different regions along the plate. For the DC voltage, the left (blue) and right (red) electrodes correspond to negative and positive charges, respectively. The order parameter profile shows a peak on top of both electrodes, indicating that the maximum deviation from the initial ordering occurs near the electrode positions. Interestingly, the peaks are symmetrical for AC potential and asymmetrical with a large peak on the negative electrode for the DC potential. With this, we can obtain a complete picture of the water molecule ordering near the electrode surface. In the absence of an applied electric field, water molecules at the gold interface exhibit preferential orientation driven by subtle van der Waals and image charge interactions. Studies have shown that the molecular dipole of water tends to lie parallel to the gold surface, resulting in an in-plane orientation and a negative order parameter ($C_{z}<0$), as also observed in the present study [limmer2011putative, bonthuis2011dielectric]. In this configuration, the oxygen atom (O) typically lies closer to the gold surface, while the hydrogen atoms (H) point slightly away from the surface. This geometry arises because oxygen interacts more favorably with the surface, potentially through weak hydrogen bonding or van der Waals attraction, while the hydrogen atoms avoid direct interaction with gold due to their weaker or slightly repulsive interactions. As a result, the dipole vector (from O to the midpoint of the H–H bond) lies parallel to the surface or is slightly tilted away, leading to a well-defined in-plane ordering in the interfacial water layer.
@@
When an electric potential is applied, the orientation of these dipoles changes in response to the polarity of the electrodes. At the negative electrode, the positively charged hydrogen atoms are attracted towards the surface. This causes the dipoles to tilt in the opposite direction compared to their initial configuration, disrupting the original in-plane alignment. As a result, the order parameter $C_{z}$ increases significantly towards zero, indicating a strong reorientation or near-random dipole alignment. This explains why, under DC potential, the $C_{z}$ peak on top of the negative electrode rises to nearly zero.
@@
In contrast, at the positive electrode, the oxygen atoms, which were already slightly tilted towards the surface, experience further attraction due to their negative partial charge. This reinforces the initial orientation rather than disrupting it, resulting in only a modest change in the order parameter. Consequently, the $C_{z}$ peak on the positive side remains closer to 0.18-0.18, not reaching zero. This asymmetry in reorientation between the two electrodes under DC voltage accounts for the asymmetric peak heights in the $C_{z}$ profile observed in Fig. 9. For the AC case, the field reverses direction periodically, which averages out these polarity-specific effects, leading to a more symmetrical response in the order parameter profile.
@@
Until now, we have been focussing on symmetrical electrodes with the same dimensions. It has been already reported that there is a net flow of electrolyte solutions when AC potential is applied to asymmetric pairs of the electrode structure [ramos2003pumping, ajdari1996generation]. Here we focus on three different models to study the influence of asymmetry of electrodes. The first electrode structure is a symmetric pair of electrodes with both electrodes of the same length D and the distance between the electrodes is also fixed at D. The second structure is an asymmetric electrode in which the left electrode length is D and the right electrode length is 2D (twice the length of left electrode) and the distance between the electrodes is D. Similarly, the third electrode structure is an asymmetric electrode in which the left electrode length is D and the right electrode is 2D, however the distance between electrodes is 2D. These models are named as model 1, model 2, and model 3 respectively. The structure and dimensions of the three models are shown in Fig. 10. MD simulations were carried out on these three models using the CPM on the electrodes with a 4 V AC potential having a frequency of 100 GHz. The $x$ component of the fluid velocity in the $x$-$y$ plane for the three models is plotted in Fig. 10. The velocity profile is averaged over five different independent simulations. Due to the periodicity of the system, even in the cases where the electrodes have different lengths there is symmetry (through reflection + translation). This means that if flow is produced by the AC voltage, it is equally likely to be in either direction. Therefore, in averaging the five independent simulations, the average velocities reported are relative to the sign of the overall velocity of the first simulation.
The average value of the $x$ velocity component in the entire channel is given in the legend of the plots of Fig. 10. From the velocity profile, it is evident that a net flow in the $x$ direction is obtained for asymmetric electrode configuration. Whereas the net velocity of the symmetric electrode structure is zero, within the statistical uncertainty. It is also noted that model 3 gives maximum velocity. The average value of the $x$ component of model 3 is 0.78 $\pm$ 0.34 m s-1. A parabolic flow with maximum velocity at the center of the channel is obtained. Similarly, model 2 also shows a net $x$ component velocity of 0.23 $\pm$ m 0.09s-1.
While these velocities are macroscopically significant, the resulting translational kinetic energy per mole of water molecules (5.5 mJ mo-1 at speeds of 0.78 m s-1) is still six orders of magnitude below the thermal energy scale ($RT=2.5$ kJ mol-1 when $T=300$ K and R is the ideal gas constant).
@@
Figure 10 also shows the streamline plot for the three models. The streamlines are more random and chaotic in model 1. The absence of a significant net flow in any direction is also confirmed by the average velocity profile which indicates no net flow. This implies that the forces acting on the system are balanced. It results in a homogeneous distribution of the streamlines with no dominance of the directional movement.
@@
The third model displays highly directional streamlines, thereby indicating a strong and coherent net flow. This suggests that the forces driving the flow are strong and dominate over turbulent influences. The mean velocity profile shows a net flow in the $x$-direction having a velocity component of 0.78 $\pm$ 0.34 m s-1. This net flow in the $x$ direction, together with the streamlined directionality pattern, suggests there is a preferential flow direction. The asymmetries in forces acting on the system drive the fluid preferentially along the $x$ axis over the period considered.
@@
The mean square displacement (MSD) of the molecules was calculated to understand the fluid properties, and the results are shown in Fig. 11. These systems have been analyzed under different electric field conditions: 0.0 V (no field), 4.0 V DC , and 4.0 V at 100 GHz AC. The MSD plots help to depict how the position of the molecules changes with time. At 0.0 V, a linear increase in MSD indicates normal diffusive behavior. At 4.0 V DC, the MSD is close to that of the no-field condition, indicating minimal impact on the mobility. However, the 100 GHz AC electric field of 4.0 V results in a significantly increased MSD, indicating that the mobility molecules is being enhanced. The increase in MSD observed under 100 GHz AC excitation arises indirectly from the temperature rise induced by the high-frequency electric field. As shown in 3, no comparable temperature increase or MSD enhancement is observed under 0 V or DC conditions, indicating that the enhanced diffusion is driven by AC-field-induced heating rather than by a direct, non-thermal effect of the electric field on molecular mobility
@@
Our study aimed to uncover the molecular-level mechanisms responsible for the AC-EHD flow. Several explanations regarding the underlying mechanism of the AC-EHD flow have been proposed in the literature [ramos1998ac, jones1995electromechanics]. Here, by combining the results of our MD simulation and previous understandings, we provide further insight into the mechanism. Various important factors are responsible for this phenomenon. First, localized heat generation occurs near the electrodes due to the effect of the rapid reorientation of water dipoles under the high-frequency AC potential [incropera1996fundamentals]. This localized heating gives rise to a steep temperature gradient between the region near the electrodes and the surrounding cooler fluid. At nanometer length scales, such temperature gradients modify local fluid properties and the response of the fluid to the applied electric field. Under asymmetric conditions, this leads to field-induced fluid motion that enhances transport and mixing within the channel [white2011fluid].
Secondly, electrothermal effects also play a significant role. The electric field changes the temperature-dependent properties of water, such as its conductivity and permittivity, to give rise to these effects [ramos1998ac]. This gradient in conductivity and permittivity, in turn, interacts with the electric field to produce electrothermal forces that drive the fluid movement. The electrothermal forces increase at higher frequencies, thereby enhancing the fluid movement [castellanos2014electrohydrodynamics].
@@
Nevertheless, a net flow is only visible when the electrodes are asymmetric. In symmetric electrode configurations, the aforementioned forces are balanced, thereby eliminating the chances of a net directional flow. The difference in the electrode dimensions and the distance between electrodes will enhance the imbalance in forces acting on the fluid, leading to a pronounced AC-EHD effect. Thus, the high-frequency AC potential enhances molecular mobility while inducing localized heating near the electrodes. The combined influence of field-induced thermal effects and non-uniform electric fields drives fluid motion under AC-EHD conditions. Notably, asymmetry in the electrode configuration is essential for creating an imbalance in these effects, resulting in a net directional flow. The present results provide qualitative guidance for the design of nanoscale AC-EHD devices. The simulations indicate that operation at very high AC frequencies is essential to activate ion-independent flow under strong confinement, while geometric asymmetry in electrode configurations is a necessary condition for generating directed motion. In addition, the weak dependence of the flow on ionic concentration suggests potential advantages for applications involving complex or variable electrolytes, such as biosensing and lab-on-a-chip systems. The highly localized nature of the flow and heating near electrode surfaces further suggests opportunities for targeted transport or mixing close to functional interfaces rather than bulk stirring.
@@ IV CONCLUSIONS
Several physical mechanisms were identified as contributors to this flow: (1) field-induced effects associated with temperature-induced density gradients, (2) electrothermal forces from the interaction between the electric field and temperature-dependent properties of water, and (3) Maxwell stress2) stresses arising from non-uniform electric fields and electrostriction caused by spatially varying polarization. The observed flow is enabledemerges only when these forceseffects are imbalanced, which occurs under geometrically asymmetric electrode configurations.
@@ Acknowledgements.
This research was supported by the Australian Research Council through the ARC Laureate Fellowship program (FL190100080 and FL220100059). It was also supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia and resources and services from the National Computational Infrastructure (NCI) under the Adapter scheme, which is supported by the Australian Government. We also wish to acknowledge The University of Queensland’s Research Computing Centre (RCC) and Queensland Cyber Infrastructure Foundation (QCIF) for their support.